# Experimental Evidence for the Propagation and Preservation of Machine Discoveries in Human Populations

### v1

@@ Abstract
Machines powered by Artificial Intelligence (AI) have the potential to discover strategies that are difficult for humans to conceive: their computational capacity exceeds human processing power and speed; their ability to process and share information in serial or parallel with high fidelity allows them to distribute problem-solving efficiently [1]. AlphaGo, an AI system that defeated world champion Lee Sedol in the ancient game of Go, discovered superhuman strategies by playing millions of games against a copy of itself, and integrated these experiences into a complex strategy encoded in its neural network [2]. Given the human-defined objective of winning the game, the algorithm seized extensive search to identity a successful strategy that has been described as ”alien” considering thoughts of year of human Go-play[3]. Recent evidence suggests that humans are beginning to change their own Go-play as a result of this innovation [4]. The domain of gameplay remains, however, a rare example of human problem-solving strategies being shaped by machine-discovered strategies. This raises a question: Beyond being mere tools, under which conditions can machines persistently reshape the way we solve problems?
We address this question by integrating insights from the field of Cultural Evolution [5]: culture changes through a process of variation, transmission, and selection [6]. We propose that, correspondingly, machine-induced cultural shifts are contingent on the convergence of three critical conditions. Appropriate Discovery Difficulty: Strategies must be non-trivial and extend beyond the range of variations typically discovered by humans. Low Transmission Difficulty: Humans must be able to learn and effectively disseminate machine-introduced discoveries. Recognizable Selective Advantage: Cultural adoption of machine discoveries hinges on their demonstrated success and the extent to which humans attend to them and endorse them. As such, our proposal integrates strands of literature that have remained largely separate: prior research highlighting the transformative power of selective social learning [7, 8, 9]; observational evidence for the general possibility of machine-induced cultural shifts [2, 4]; and conjectures about the complementary potential of human limitations and machine capacities [1].
@@
Even though intelligent machines are increasingly used in domains of human discovery, such as education and research [22, 23, 24], in most cases these AI systems do not convey problem-solving strategies that are entirely alien to human culture. As such, they do not satisfy the above conditions but remain within the confines of existing human discoveries (see Figure 1B bottom left). And even those AI systems that have discovered super-human knowledge, remain –at present– tools in human hands rather than integrated into human cognitive frameworks. AlphaFold provides a remarkable example for this category: while it was inspired by the success of AlphaGo in emulating the intuition of human experts [25] and in doing so has achieved solving protein-folding problems at superhuman levels [26, 27], its discoveries have not (yet) transpired into novel human intuitions, unlike in the case of AlphaGo. As such, it presents an unlearnable and unpreservable machine discovery (see Figure 1B top right).
@@
To provide the first experimental evidence of preserved, machine-induced cultural shift in humans, we designed an experimental setup that features conditions of appropriate discovery and transmission difficulty, along with a recognizable selective advantage for machine performance. The task emulated a prototypical situation in which human exploration is constrained by the pervasive human tendency to avoid losses (i.e., myopic decision-making). Human participants (n = 1,155) were paid to solve a series of “reward networks” (Figure 2 C), which we developed as a modified version of an established paradigm [28, 29]. In our task, networks consisted of nodes connected by links that were associated with rewards ranging from -50 to +400 points. Participants’ goal was to choose 10 moves through the network that would maximize the number of points gained. Our networks were designed such that there was an optimal strategy for solving them: incurring early losses subsequently led to links associated with the maximum reward of +400 points, which were not accessible when avoiding early losses. Hence, the optimal strategy for solving our networks contradicts the myopic tendency of humans to avoid losses and prioritize immediate gains, even when accepting short-term losses could be beneficial in the long run. In line with prior literature, we expected that this myopic tendency would make the discovery of optimal strategy in our task difficult for human players [30]. Meanwhile, we expected that a few minutes of training on our task would be sufficient for a deep-Q-learning algorithm (the ”machine player”) to successfully discover the optimal strategy. To compare the performance of human players on their own against the performance of human players exposed to machine players, we randomly allocated participants into populations that involved either only eight humans (15 ”Human-Only Populations”) or five humans and three machine players in their initial generation (15 ”Human-Machine Populations”). Each population consisted of five generations (numbered 0-4; Fig. 2 A), allowing us to observe differences between the two kinds of populations over time. After a series of individual attempts at solving networks, participants in Generation 0 solved four networks during the ”Demonstration Phase”, knowing that their moves would be recorded for future generations to learn from. Participants from Generations 1-4 could choose a ”demonstrator” from the previous generation to learn from, based on the demonstrators’ average performance. They did not know the demonstrators’ identity as a human or machine, but were informed that machine players were a possibility. Only after this Social Learning Phase did participants from Generations 1-4 demonstrate their moves for the next generation. All networks solved by the participants throughout the experiment were unique, requiring them to discover or learn a generalizable strategy rather than optimizing a specific solution.
@@
Panel A in Figure 4 shows the average reward gained in the task by generation and population. Descriptively, while the two types of populations started out with the same average in Generation 0, in Generations 1-4 Human-Machine Populations consistently earned higher rewards. Our preregistered linear mixed-effects model on these latter generations supported that this difference was significant ($\beta=-309.2$, SE = $57.0$, $\text{CI}_{95\%}$ = [$-420.7$, $-196.6$]). Further models confirmed that Human-Machine Populations performed better than Human-Only Populations both in Generation 1, which could learn directly from machine players ($\beta=-404.6$, SE = $59.4$, $\text{CI}_{95\%}$ = [$-513.5$, $-293.0$]), and in Generation 4, where influence from machines was remote by four iterations ($\beta=-229.1$, SE = $82.4$, $\text{CI}_{95\%}$ = [$-407.4$, $-72.3$]).
@@
Panel B in Figure 4 shows the proportion of human moves congruent with machine-generated moves, summarized by generation and population. Qualitatively, the results are similar to the results for rewards: no difference between conditions in Generation 0, but Human-Machine Populations are consistently more behaviourally congruent in Generations 1 to 4. A logistic mixed-effects model on these generations shows that Human-Machine Populations were more congruent than human populations ($\beta=-0.88$, SE = $0.14$, $\text{CI}_{95\%}$ = [$-1.14$, $-0.54$]). Again, this was also the case within Generation 1, which learned from machine players ($\beta=-1.16$, SE = $0.14$, $\text{CI}_{95\%}$ = [$-1.43$, $-0.90$]), and Generation 4 ($\beta=-0.66$, SE = $0.19$, $\text{CI}_{95\%}$ = [$-1.03$, $-0.24$]).
@@
During the Demonstration Phase of our experiment, players were asked to provide a written strategy of how they approached the task. In contrast to the demonstration itself, these written strategies were not transmitted to the next generation, and players were aware of this. Panel C in Figure 4 shows the proportion of human players mentioning the optimal strategy (i.e. the ”loss strategy”) in their written statements, as coded independently by three human annotators. In the Human-Only Populations, we observe practically no mention of the strategy, with proportions close to 0 throughout all generations; the exception to this is the single population that discovered the optimal strategy in Generation 3, resulting in 50% of players referring to it in Generation 4. In contrast, participants in the Human-Machine Populations consistently mentioned the optimal strategy explicitly, with proportions of about 30% to 50% per generation. A logistic mixed-effects model on Generations 1 to 4 shows that this difference is significant ($\beta=-4.18$, SE = $0.62$, $\text{CI}_{95\%}$ = [$-5.49$, $-3.04$]).
@@
We devised an agent-based simulation to systematically explore how variations in discovery and transmission difficulty shape the adoption of machine-discoveries and, ultimately, population-level success. Corroborating our experimental results, harder-to-discover strategies (with a higher (log(d)) were adopted when transmission difficulty $1-t$ was low, where $t$ is the transmission rate towards the end of the multi-generational simulation (Panel A in Fig. 5). This effect was enhanced when agents selected the highest-performing individual from the previous generation to socially learn from (solid lines) compared to learning from a random agent (dashed lines). When machine agents, with their increased discovery rate, were added, these human-machine populations could discover and preserve the optimal solution even when discovery was difficult (orange solid line). This machine-induced effect was not present when agents learned from a random agent of the previous generation (orange dashed line).
@@ Discussion
Acknowledgments
Author contributions:
Competing interests:
There are no competing interests to declare.
Data and materials availability:
Supplementary materials
Materials and Methods
Supplementary Text
Figs. S1 to S10
Tables S1 to S3
Supplementary Materials for
Levin Brinkmann, Thomas F. Eisenmann∗†, Anne-Marie Nussberger,
Maxime Derex, Sara Bonati, Valerii Chirkov, Iyad Rahwan
Corresponding authors. Email: mueller@mpib-berlin.mpg.de (T.F.E.); rahwan@mpib-berlin.mpg.de (I.R.)
These authors contributed equally to this work.
This PDF file includes:
Materials and Methods
Supplementary Text
Figures S1 to S10
Tables S1 to S3
Materials and Methods
@@ Acknowledgments
https://center-for-humans-and-machines.github.io/reward-network-iii.
@@
We conducted the simulation for two population configurations: one consisting solely of human agents and another including three machine agents in the first generation. The simulations were run across a grid of discovery difficulties ($-\log{d_{\text{optimal}}}$) and transmission difficulties ($1-t$) with 100 replications for each parameter combination. As a robustness check, Supplementary Figure S11 includes results for simulations with a single machine in generation 0, participant selection from all previous-generation agents, and an increased population size of 16.
@@
We ran the linear mixed-effects model testing prediction 1a (”General Benefit of AI”) on the demonstration trials of generations 1 to 4, with total reward as the dependent variable and fixed effects for condition, generation, and their interaction. Additionally, we included random intercepts for participant and population, as well as a random slope for generation on population. As preregistered, we centered and scaled generation to improve model convergence and make the main effect for condition interpretable. In line with our prediction, the model revealed a significant main effect for condition ($\beta=-309.2$, SE = $57.0$, $\text{CI}_{95\%}$ = [$-418.4$, $-200.1$]), indicating that human-machine populations outperformed human populations overall.
The model for prediction 1b (”Long-lasting Benefit of AI”) followed a similar logic, but was using only the data of generation 4. Given the more restricted data set, only the fixed effect for condition and the random intercepts for participant and population remained. Similar to the first model, there was a meaningful main effect for condition ($\beta=-229.1$, SE = $82.4$, $\text{CI}_{95\%}$ = [$-383.1$, $-55.5$]), supporting our prediction that human-machine populations would perform better than human populations by the last generation.
The model for prediction 2a (”Transmission from AI to Humans”) was identical to the one for 1b, but used the data of generation 1. Again, there was a significant main effect for condition ($\beta=-404.6$, SE = $59.4$, $\text{CI}_{95\%}$ = [$-524.8$, $-286.1$]), supporting our prediction that human-machine populations performed better than human populations in the immediate generation learning from machine agents. Here, the model returned a warning about a singular fit, so we also checked all other optimizers available in lme4; however, since all optimizers converged to identical values for the fixed effects, we concluded the warning to be false-positive.
@@
The models assessing predictions 1a to 2a were identical to the ones for reward, with exception of the different outcome measure and being on the level of moves per trial. Since behavioural congruency was a binary variable, we used a logit link here. All three models replicated the expected effects for condition found for reward on the behavioural congruency measure: Human-machine populations were behaviourally more congruent than human populations overall ($\beta=-0.88$, SE = $0.14$, $\text{CI}_{95\%}$ = [$-1.16$, $-0.59$]), in the last generation ($\beta=-0.66$, SE = $0.19$, $\text{CI}_{95\%}$ = [$-1.02$, $-0.31$]), and in the first generation ($\beta=-1.16$, SE = $0.14$, $\text{CI}_{95\%}$ = [$-1.42$, $-0.90$]). Again, the last model returned a warning about a singular fit, but all available optimizers converged to identical fixed effects until the second digit, so we took it as a false positive.
@@
The model assessing prediction 2b (”Explicit Recognition of Counterintuitive Strategy”) was identical to the one for prediction 1a, but using a logit link due to the binary outcome variable and missing the random intercept for participants, since it was built on participant level. Again, there was a significant main effect for condition ($\beta=-4.18$, SE = $0.62$, $\text{CI}_{95\%}$ = [$-6.30$, $-3.11$]), indicating that human-machine populations explicitly mentioned the loss strategy more frequently than human populations from generations 1 to 4. This model also returned a warning about a singular fit, but all available optimizers converged to qualitatively similar parameters.
@@
Condition = Human Population --309.2 --229.1 --404.6
@@
Condition = Human Population --0.884 --0.657 --1.159
@@
Condition = Human Population --4.180
[$-6.30$, $-3.11$]
Generation --0.503
[$-0.89$, $-0.10$]

### v2

@@ Abstract
Machines powered by Artificial Intelligence (AI) have the potential to discover strategies that are difficult for humans to conceive: their computational capacity exceeds human processing power and speed; their ability to process and share information in serial or parallel with high fidelity allows them to distribute problem-solving efficiently [1]. AlphaGo, an AI system that defeated world champion Lee Sedol in the ancient game of Go, discovered superhuman strategies by playing millions of games against a copy of itself, and integrated these experiences into a complex strategy encoded in its neural network [2]. Given the human-defined objective of winning the game, the algorithm used extensive search to identify a successful strategy that has been described as ”alien”, considering thousands of years of human Go-play [3]. Recent evidence suggests that humans have significantly improved their own Go-play as a result of this innovation [4], while more subtle shifts that appear less ”alien” when viewed in historical context are also evident in Go opening moves [5]. In chess, selected strategic concepts extracted from an AI system were applied by grandmasters after observing the AI system only on a few examples [6]. The domain of gameplay remains, however, a rare example of human problem-solving strategies being shaped by machine-discovered strategies. This raises a question: Beyond being mere tools, under which conditions can machines persistently reshape the way we solve problems?
We address this question by integrating insights from the field of Cultural Evolution [7]: culture changes through a process of variation, transmission, and selection [8]. We propose that, correspondingly, machine-induced cultural shifts are contingent on the convergence of three critical conditions. Appropriate Discovery Difficulty: Strategies must be non-trivial and extend beyond the range of variations typically discovered by humans. Low Transmission Difficulty: Humans must be able to learn and effectively disseminate machine-introduced discoveries. Recognizable Selective Advantage: Cultural adoption of machine discoveries hinges on their demonstrated success and the extent to which humans attend to them and endorse them. As such, our proposal integrates strands of literature that have remained largely separate: prior research highlighting the transformative power of selective social learning [9, 10, 11]; observational evidence for the general possibility of machine-induced cultural shifts [2, 4, 6]; and conjectures about the complementary potential of human limitations and machine capacities [1].
@@
Even though intelligent machines are increasingly used in domains of human discovery, such as education and research [24, 25, 26], in most cases these AI systems do not convey problem-solving strategies that are entirely alien to human culture. As such, they do not satisfy the above conditions but remain within the confines of existing human discoveries (see Figure 1B bottom left). And even those AI systems that have discovered super-human knowledge, often remain – at present – tools in human hands rather than integrated into human cognitive frameworks. AlphaFold provides a remarkable example for this category: while it was inspired by the success of AlphaGo in emulating the intuition of human experts [27] and in doing so has achieved solving protein-folding problems at superhuman levels [28, 29], its discoveries have not (yet) transpired into novel human intuitions, unlike in the case of AlphaGo and AlphaZero. As such, it presents an unlearnable and unpreservable machine discovery (see Figure 1B top right).
@@
To provide the first experimental evidence of preserved, machine-induced cultural shift in humans, we designed a behavioral task that features conditions of appropriate discovery and transmission difficulty, along with a recognizable selective advantage for machine performance. The task emulated a prototypical situation in which human exploration is constrained by the pervasive human tendency to avoid losses (i.e., myopic decision-making). Human participants (n = 1,155) were paid to solve a series of “reward networks” (Figure 2 C), which we developed as a modified version of an established paradigm [30, 31]. In our task, networks consisted of nodes connected by links that were associated with rewards ranging from -50 to +400 points. Participants’ goal was to choose 10 moves through the network that would maximize the number of points gained. Our networks were designed such that there was an optimal strategy for solving them: incurring early losses subsequently led to links associated with the maximum reward of +400 points, which were not accessible when avoiding early losses. Hence, the optimal strategy for solving our networks contradicts the myopic tendency of humans to avoid losses and prioritize immediate gains, even when accepting short-term losses could be beneficial in the long run. In line with prior literature, we expected that this myopic tendency would make the discovery of optimal strategy in our task difficult for human players [32]. Meanwhile, we expected that a few minutes of training on our task would be sufficient for a deep-Q-learning algorithm (the ”machine player”) to successfully discover the optimal strategy. To compare the performance of human players on their own against the performance of human players exposed to machine players, we randomly allocated participants into populations that involved either only eight humans (15 ”Human-Only Populations”) or five humans and three machine players in their initial generation (15 ”Human-Machine Populations”). Each population consisted of five generations (numbered 0-4; Fig. 2 A), allowing us to observe differences between the two kinds of populations over time. After a series of individual attempts at solving networks, participants in Generation 0 solved four networks during the ”Demonstration Phase”, knowing that their moves would be recorded for future generations to learn from. Participants from Generations 1-4 could choose a ”demonstrator” from the previous generation to learn from, based on the demonstrators’ average performance. They did not know the demonstrators’ identity as a human or machine, but were informed that machine players were a possibility. Only after this Social Learning Phase did participants from Generations 1-4 demonstrate their moves for the next generation. All networks solved by the participants throughout the experiment were unique, requiring them to discover or learn a generalizable strategy rather than optimizing a specific solution.
@@
Panel A in Figure 4 shows the average reward gained in the task by generation and population. Descriptively, while the two types of populations started out with the same average in Generation 0, in Generations 1-4 Human-Machine Populations consistently earned higher rewards. Our preregistered linear mixed-effects model on these latter generations supported that this difference was significant ($\beta=-309.2$, SE = $57.0$, $\text{CI}_{95\%}$ = [420.7-420.7, 196.6-196.6]). Further models confirmed that Human-Machine Populations performed better than Human-Only Populations both in Generation 1, which could learn directly from machine players ($\beta=-404.6$, SE = $59.4$, $\text{CI}_{95\%}$ = [513.5-513.5, 293.0-293.0]), and in Generation 4, where influence from machines was remote by four iterations ($\beta=-229.1$, SE = $82.4$, $\text{CI}_{95\%}$ = [407.4-407.4, 72.3-72.3]).
@@
Panel B in Figure 4 shows the proportion of human moves congruent with machine-generated moves, summarized by generation and population. Qualitatively, the results are similar to the results for rewards: no difference between conditions in Generation 0, but Human-Machine Populations are consistently more behaviourally congruent in Generations 1 to 4. A logistic mixed-effects model on these generations shows that Human-Machine Populations were more congruent than human populations ($\beta=-0.88$, SE = $0.14$, $\text{CI}_{95\%}$ = [1.14-1.14, 0.54-0.54]). Again, this was also the case within Generation 1, which learned from machine players ($\beta=-1.16$, SE = $0.14$, $\text{CI}_{95\%}$ = [1.43-1.43, 0.90-0.90]), and Generation 4 ($\beta=-0.66$, SE = $0.19$, $\text{CI}_{95\%}$ = [1.03-1.03, 0.24-0.24]).
@@
During the Demonstration Phase of our experiment, players were asked to provide a written strategy of how they approached the task. In contrast to the demonstration itself, these written strategies were not transmitted to the next generation, and players were aware of this. Panel C in Figure 4 shows the proportion of human players mentioning the optimal strategy (i.e. the ”loss strategy”) in their written statements, as coded independently by three human annotators. In the Human-Only Populations, we observe practically no mention of the strategy, with proportions close to 0 throughout all generations; the exception to this is the single population that discovered the optimal strategy in Generation 3, resulting in 50% of players referring to it in Generation 4. In contrast, participants in the Human-Machine Populations consistently mentioned the optimal strategy explicitly, with proportions of about 30% to 50% per generation. A logistic mixed-effects model on Generations 1 to 4 shows that this difference is significant ($\beta=-4.18$, SE = $0.62$, $\text{CI}_{95\%}$ = [5.49-5.49, 3.04-3.04]).
@@
We devised an agent-based simulation to systematically explore how variations in discovery and transmission difficulty shape the adoption of machine-discoveries and, ultimately, population-level success. Corroborating our experimental results, harder-to-discover strategies (with a higher (log(d)) were adopted when transmission difficulty 1t1-t was low, where $t$ is the transmission rate towards the end of the multi-generational simulation (Panel A in Fig. 5). This effect was enhanced when agents selected the highest-performing individual from the previous generation to socially learn from (solid lines) compared to learning from a random agent (dashed lines). When machine agents, with their increased discovery rate, were added, these human-machine populations could discover and preserve the optimal solution even when discovery was difficult (orange solid line). This machine-induced effect was not present when agents learned from a random agent of the previous generation (orange dashed line).
@@ Discussion
Methods
@@
We conducted the simulation for two population configurations: one consisting solely of human agents and another including three machine agents in the first generation. The simulations were run across a grid of discovery difficulties ($\log{d_{\text{optimal}}}$) and transmission difficulties (1t1-t) with 100 replications for each parameter combination. As a robustness check, Supplementary Figure S11 includes results for simulations with a single machine in generation 0, participant selection from all previous-generation agents, and an increased population size of 16.
Acknowledgments
Author contributions:
Competing interests:
There are no competing interests to declare.
Data and materials availability:
Supplementary information
Supplementary Text
Figs. S1 to S10
Tables S1 to S3
Supplementary Information for
Levin Brinkmann, Thomas F. Eisenmann∗†, Anne-Marie Nussberger,
Maxime Derex, Sara Bonati, Valerii Chirkov, Iyad Rahwan
Corresponding authors. Email: eisenmann@mpib-berlin.mpg.de (T.F.E.); rahwan@mpib-berlin.mpg.de (I.R.)
These authors contributed equally to this work.
This PDF file includes:
Supplementary Text
Figures S1 to S10
Tables S1 to S3
@@ Acknowledgments
We ran the linear mixed-effects model testing prediction 1a (”General Benefit of AI”) on the demonstration trials of generations 1 to 4, with total reward as the dependent variable and fixed effects for condition, generation, and their interaction. Additionally, we included random intercepts for participant and population, as well as a random slope for generation on population. As preregistered, we centered and scaled generation to improve model convergence and make the main effect for condition interpretable. In line with our prediction, the model revealed a significant main effect for condition ($\beta=-309.2$, SE = $57.0$, $\text{CI}_{95\%}$ = [418.4-418.4, 200.1-200.1]), indicating that human-machine populations outperformed human populations overall.
The model for prediction 1b (”Long-lasting Benefit of AI”) followed a similar logic, but was using only the data of generation 4. Given the more restricted data set, only the fixed effect for condition and the random intercepts for participant and population remained. Similar to the first model, there was a meaningful main effect for condition ($\beta=-229.1$, SE = $82.4$, $\text{CI}_{95\%}$ = [383.1-383.1, 55.5-55.5]), supporting our prediction that human-machine populations would perform better than human populations by the last generation.
The model for prediction 2a (”Transmission from AI to Humans”) was identical to the one for 1b, but used the data of generation 1. Again, there was a significant main effect for condition ($\beta=-404.6$, SE = $59.4$, $\text{CI}_{95\%}$ = [524.8-524.8, 286.1-286.1]), supporting our prediction that human-machine populations performed better than human populations in the immediate generation learning from machine agents. Here, the model returned a warning about a singular fit, so we also checked all other optimizers available in lme4; however, since all optimizers converged to identical values for the fixed effects, we concluded the warning to be false-positive.
@@
The models assessing predictions 1a to 2a were identical to the ones for reward, with exception of the different outcome measure and being on the level of moves per trial. Since behavioural congruency was a binary variable, we used a logit link here. All three models replicated the expected effects for condition found for reward on the behavioural congruency measure: Human-machine populations were behaviourally more congruent than human populations overall ($\beta=-0.88$, SE = $0.14$, $\text{CI}_{95\%}$ = [1.16-1.16, 0.59-0.59]), in the last generation ($\beta=-0.66$, SE = $0.19$, $\text{CI}_{95\%}$ = [1.02-1.02, 0.31-0.31]), and in the first generation ($\beta=-1.16$, SE = $0.14$, $\text{CI}_{95\%}$ = [1.42-1.42, 0.90-0.90]). Again, the last model returned a warning about a singular fit, but all available optimizers converged to identical fixed effects until the second digit, so we took it as a false positive.
@@
The model assessing prediction 2b (”Explicit Recognition of Counterintuitive Strategy”) was identical to the one for prediction 1a, but using a logit link due to the binary outcome variable and missing the random intercept for participants, since it was built on participant level. Again, there was a significant main effect for condition ($\beta=-4.18$, SE = $0.62$, $\text{CI}_{95\%}$ = [6.30-6.30, 3.11-3.11]), indicating that human-machine populations explicitly mentioned the loss strategy more frequently than human populations from generations 1 to 4. This model also returned a warning about a singular fit, but all available optimizers converged to qualitatively similar parameters.
@@
Condition = Human Population -309.2 -229.1 -404.6
@@
Condition = Human Population -0.884 -0.657 -1.159
@@
Condition = Human Population -4.180
[6.30-6.30, 3.11-3.11]
Generation -0.503
[0.89-0.89, 0.10-0.10]